OpenStudy (anonymous):

A firecracker shoots up from a hill 140 feet high with an initial speed of 100 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? 5 seconds 7 seconds 9 seconds 11 seconds Question 18 (Multiple Choice Worth 5 points) [9.06] Jacob kicks a soccer ball off the ground and in the air with an initial velocity of 33 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 15.1 feet 16.5 feet 17.0 feet 18.2 feet

OpenStudy (anonymous):

\[H(t) = −16t^2 + vt + s\] \[v=100,s=140\implies H(t) = −16t^2 + 100t + 140\]

OpenStudy (anonymous):

set \[-16t^2+100t+140=0\] and solve for \(t\) which will tell you how long before it hits the ground (since when it hits the ground the height is 0)

OpenStudy (anonymous):

dont get it

OpenStudy (anonymous):

do you know what \[H(t) = −16t^2 + vt + s\] means?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then of course you do not get it! it is supposed to represent the height of the firecracker in feet, if \(t\) is time in seconds. the initial velocity is \(v\) and you are told that \(v=100\) the initial height is \(s\) and you are told \(s=140\) you get \[H(t) = −16t^2 + 100t + 140\] so for example after 1 second it will be \[H(1) = −16\times 1^2 + 100\times 1 + 140=224\] i.e. after one second it will be 224 feet in the air

OpenStudy (anonymous):

oh got it

OpenStudy (anonymous):

whats the answer